PRACTICE IT Use the worked example above to help you solve this problem. A ball is thrown upward from the top of a building at an angle of 30.00 to the horizontal and with an initial speed of 19.0 m/s. The point of release is h 46.0 m above the ground. (a) How long does it take for the ball to hit the ground? 4.18 b) Find the balls speed at impact 35.5 m/s (c) Find the horizontal range of the ball. 68.78 EXERCISE HINTS: GETTING STARTED I IM STUCK! Suppose the ball is thrown from the same height as in the PRACTICE IT problem at an angle of 39.0 below the horizontal. If it strikes the ground 35.9 m away, find the following. (Hint: For part (a), use the equation for the x-displacement to eliminate vot from the equation for the y-displacement.) (a) the time of flight (b) the initial speed m/s (c) the speed and angle of the velocity vector with respect to the horizontal at impact speed angle m/s o below the horizontal Need Help? Talk to Tutor Submit Answer Save Progress
Answer #1

35. 1 test y.Il co, 29 6 111-22 mis

EDIT:

(a) There was math error.

Correct answer is: t = 1.86 s

(b) v = 24.9 m/s

(c) Vx = v cos 39 = 19.34 m/s

Vy = v sin39 + gt = 33.9 m/s

So, v = sqrt (Vx^2 + Vy^2) = 39.0 m/s

Dir =arctan (Vy/Vx) = 60.3 degrees